OpenStudy (anonymous):

. Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8.

OpenStudy (anonymous):

A. y=32x B. Y^2= 8x C. Y= 1/32 x^2 D. y= 1/8 x^2

OpenStudy (jdoe0001):

so... hmmm the focus is at (0,8) an the directrix is y = -8 can you see where the vertex is at?

OpenStudy (jdoe0001):

keep in mind, the vertex is half-way between both of those fellows the vertex is half-way between the focus point and the directrix

OpenStudy (anonymous):

so wouldn't it be choice B.? because of the Y

OpenStudy (jdoe0001):

well... we don't have a vertex really... so... we dunnno what it might be

OpenStudy (anonymous):

oh okasy would you mind showing me how to do the problem im not sure how to di it!!! @jdoe0001

OpenStudy (jdoe0001):